1.A fixed mass of gas has 248cm³ at temperature of _20degress and 760mmHg pressure. Calculate the temperature in degree celicius the has would occupy 458cm³ at 780mmHg pressure
2.carbon (iv)oxide takes 200s to diffuse through a porous . How long will it take for the same amount of carbon (iv)oxide to diffuse through the plug under the same condition
3.A gas occupies a volume of 13dm³ at 15degress celcius and 760mm of mercury pressure.What volume would it occupy at stp and pressure
3.Thr pressure exerted by a gas in a sealed container is 1000kpa at 17degress .It was found that the container might leak if the internal pressure exceeds 1200kpa.Assuming constant volume,at what temp in degree will the container start to leak
1
Expert's answer
2020-04-28T12:26:10-0400

1) applying "\dfrac{P_1V_1}{T_1}=\dfrac{P_2V_2}{T_2}"


"\dfrac{760\times248}{253}=\dfrac{780\times458}{T_2}"

"T_2=480 K"


2) since the molar mass of both the gases are same hence the rate of diffusion will be same hence both will take same amount of time i.e. 200 seconds


3) again applying "\dfrac{P_1V_1}{T_1}=\dfrac{P_2V_2}{T_2}"


"\dfrac{13\times760}{288}=\dfrac{V_2\times760}{273}"

"V_2=12.32 dm^3"